Circular No. 6654

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cm IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (R = 20.6) on V and R CCD images taken on May 6 by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997cm is located near Abell 3528 at α = 12^{h}53^{m}37^{s}.03, δ = −28^{o}48′48′′.5 (equinox 2000.0). A nearby star has position end figures 36^{s}.51, 54′′.8. SN 1997cm was also present on similar frames taken on May 2 and Apr. 29, but it is not visible on images taken on Feb. 19 (limiting mag R = 22). Finding charts are available at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

GRB 970508 J. Heise, and J. in ’t Zand, Space Research Organization of the Nether-

lands, Utrecht; E. Costa, M. Feroci, L. Piro, P. Soffitta, L. Amati, and M. N. Cinti, Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera and G. Zavattini, Università di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; A. Tesseri, G. Gandolfi, M. Smith, H. Muller, D. Ricci, and A. Coletta, BeppoSAX, Rome; and M. Tavani, Columbia Astrophysics Laboratory and Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milan, report: “On the basis of the satellite attitude reconstruction of the BeppoSAX observation during GRB 970508, the position of the x-ray counterpart of the burst in the Wide Field Camera WFC2 is determined to be α = 6^{h}53^{m}28^{s}, δ = +79^{o}17′.4 (equinox 2000.0), with a 3′ error radius (99-percent confidence level) that is offset 2′.9 from the preliminary position reported on IAUC 6649. There are no catalogued sources within the error region. No other x-ray sources in the field-of-view of both WFCs were detected for independent verification of the satellite attitude.”

H. E. Bond, Space Telescope Science Institute, writes: “I have found a starlike variable source within the error box of GRB 970508, based on CCD observations with the Kitt Peak 0.9-m reflector. Between May 9.19 and 10.18 UT, the source brightened by ∼ 1.0 mag in the V band, and on May 10.18 was at V ∼ 20.5, V − I = +0.9. This starlike source is located at α = 6^{h}53^{m}49^{s}.2, δ = +79^{o}16′19′′ (equinox 2000.0), within both of the error boxes given above and on IAUC 6649. The variable is ∼ 3′′ east and 13′′ south of a slightly brighter star. Although the brightening (rather than fading) may make the identification with the GRB dubious, follow-up observations are desirable.”

1997 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green